OpenStudy (anonymous):

Solve the absolute value equation. │4x + 3│= 3 I forgot how to solve these o.o.....

OpenStudy (anonymous):

4x - 3 = 3 4x = 6 x = 1.5. :-)

OpenStudy (anonymous):

Oops! Oh my gosh. I saw a subtraction sign instead of addition. Just swap the signs and subtract instead of add! :(

OpenStudy (anonymous):

My final review answer key sheet says the answer is x = 0 x = - 3/2 How did they get these?

OpenStudy (anonymous):

Oh my. It seems as though I may have actually forgotten how to do these simple equations. No worries! I shall do some quick research and come back with an answer!

OpenStudy (anonymous):

Lol ok :)

OpenStudy (kc_kennylau):

|y|=a y=a or y=-a

OpenStudy (kc_kennylau):

Split |4x+3|=3 into 4x+3=3 and 4x+3=-3 :D

OpenStudy (lncognlto):

You solve it like this: │4x + 3│= 3 Therefore (4x + 3)^2 = 3^2 and so on.

OpenStudy (anonymous):

Step 1) Isolate the absolute value; 4x + 3 = 3 Step 2) Write out two equations without bars; 4x + 3 = 3 & 4x + 3 = -3! Step 3) Solve both equations; 4x + 3 = 3 ; 4x = 0 ; x = 0. 4x + 3 = -3; 4x = -6 ; x = -6 / 4 <--- Simplified = -3 / 2! :-)

OpenStudy (kc_kennylau):

@lncognlto I think that is not necessary?

OpenStudy (kc_kennylau):

@xuvaze Please do not use exclamation mark, as that confuses us.

OpenStudy (kc_kennylau):

And remember to check for extraneous (wrong) solutions after you've solved it :D

OpenStudy (lncognlto):

Possibly not, but that's how I learned to do it... I wasn't shown any other ways. xD

OpenStudy (anonymous):

I wasn't trying to be confusing. I was only expressing my passion to solve this problem. haha.

OpenStudy (kc_kennylau):

@lncognlto I'd use the definition of absolute :D |y|=a -> y=a or y=-a

OpenStudy (kc_kennylau):

@xuvaze lolz!!

OpenStudy (anonymous):

Thanks everyone :)

OpenStudy (kc_kennylau):

no problem :)

OpenStudy (anonymous):

Glad I could help! ^_^